# FGM Modeling of Thermo-Diffusive Unstable Lean Premixed Hydrogen-Air Flames

### v1

@@ 3.4 A-posteriori Unstable Flame Front Simulation
To further analyze the differences between the models, contour plots of several variables are given in Figure 8. The plots depict a zoomed section of a flame tip structure for the considered models: DC, FGM A, FGM B1, and FGM B2. The first row of images shows the normalized deviation from the unstretched mixture fraction for all models considered. The unstretched mixture fraction, $Z_{\text{1D}}(\mathcal{Y})$, is obtained from the one-dimensional flamelet, and is subtracted from the mixture fraction field to isolate the effect of stretch on the mixture fraction. The resulting value is then normalized with the unburnt mixture fraction value. The effect of preferential diffusion can be observed here: Regions with positive curvature show richer mixtures and vice versa. This results in a lean flame cusp enclosed by two rich flame bulges for DC, FGM A, and FGM B2, all of which appear very similar. In contrast, FGM B1, shows a significantly richer flame tip, which could attribute to the higher overall burning rate.
@@
In FGM B1, the enthalpy before the flame front shows an increase only just before the sharp flame tip, where curvature is highest. In areas with negative curvature, such as the cusps, there is a decrease in enthalpy instead of an increase, as shown by the DC model. This may suggest that not including the H-radical implicitly in the preferential diffusion of enthalpy limits transport along the flame front, leading to lower enthalpy values in the cusps compared to the DC model.
@@ 4 Conclusions
In this work, a planar lean premixed hydrogen flame which exhibits thermo-diffusive instabilities is simulated using the Flamelet-Generated Manifold (FGM) approach. Two manifolds are considered: manifold A, a two-dimensional manifold with progress variable and mixture fraction as control variables, and manifold B, a three-dimensional manifold with enthalpy included as an additional control variable. Choosing enthalpy as a control variable enables the inclusion of heat loss effects on the flame, which is essential for most engineering applications. From an a-priori analysis comparing manifold A and B it can be concluded that both manifolds perform well, although manifold B shows slightly improved accuracy over A in terms of relative error. The newly proposed preferential diffusion model considers only the diffusion fluxes of a limited set of major species.
A one-dimensional analysis and a linear stability analysis (LSA) suggest that for manifold A, the species H2 and H2O are sufficient, but for manifold B, the H-radical should also be included in the major species set. This is confirmed by a longer simulation in the non-linear regime. Not including the diffusion flux of the H-radical explicitly in the preferential diffusion of enthalpy leads to a poor prediction of the flame’s shape and dynamic behavior. The typical flame structures, a cusp enclosed by two bulges, observed in the DC benchmark and reported in several literature sources are not reproduced. Additionally, the flame front advances upstream much faster, which is evident from an overprediction in the burning rate. In contrast, including the H-radical diffusion flux in the preferential diffusion of enthalpy results in flame shape, dynamics, and burning rate predictions that closely match the DC benchmark. It is shown that not including the H-radical explicitly leads to a significant underprediction of the tangential enthalpy diffusion flux. The H-radical has a significant tangential component along the flame, and since it carries a significant amount of enthalpy, accurate prediction of enthalpy cross-diffusion requires the explicit inclusion of H-radical diffusion in the enthalpy preferential diffusion term.

### v2

@@ 3.4 A-posteriori Unstable Flame Front Simulation
To further analyze the differences between the models, contour plots of several variables are given in Figure 8. The plots depict a zoomed section of a flame tip structure for the considered models: DC, FGM A, FGM B1, and FGM B2. The first row of images shows the normalized deviation from the unstretched mixture fraction for all models considered. The unstretched mixture fraction, $Z_{\text{1D}}(\mathcal{Y})$, is obtained from the one-dimensional flamelet, and is subtracted from the mixture fraction field to isolate the effect of stretch on the mixture fraction. The resulting value is then normalized with the unburnt mixture fraction value. The effect of preferential diffusion can be observed here: Regions with positive curvature show richer mixtures and vice versa. This results in a lean flame cusp enclosed by two rich flame bulges for the considered models, all of which appear rather similar. The flame tip in FGM B1, however, appears slightly smaller and is enclosed by finer-scale cellular structures, which may be attributed to the shorter unstable wavelengths predicted by the linear stability analysis shown in Figure 4.
@@
In FGM B1, the enthalpy in the flame front as well as behind the flame front shows significantly lower values compared to the DC model. Furthermore, in areas with negative curvature, such as the cusps, there is a decrease in enthalpy instead of an increase, as shown by the DC model. This may suggest that not including the H-radical implicitly in the preferential diffusion of enthalpy limits transport along the flame front, leading to lower enthalpy values in the cusps compared to the DC model.
@@ 4 Conclusions
In this work, a planar lean premixed hydrogen flame which exhibits thermo-diffusive instabilities is simulated using the Flamelet-Generated Manifold (FGM) approach. Two manifolds are considered: manifold A, a two-dimensional manifold with progress variable and mixture fraction as control variables, and manifold B, a three-dimensional manifold with enthalpy included as an additional control variable. Choosing enthalpy as a control variable enables the inclusion of heat loss effects on the flame, which is essential for most engineering applications. From an a-priori analysis comparing manifold A and B it can be concluded that both manifolds perform well, although manifold B shows slightly improved accuracy over A in terms of normalized error. The newly proposed preferential diffusion model considers only the diffusion fluxes of a limited set of major species.
A one-dimensional preferential diffusion analysis and a linear stability analysis (LSA) suggest that for manifold A, the species H2 and H2O are sufficient, but for manifold B, the H-radical should also be included in the major species set. This is confirmed by a longer simulation in the non-linear regime. Not including the diffusion flux of the H-radical explicitly in the preferential diffusion of enthalpy leads to a poor prediction of the flame’s and dynamic behavior. The flame front advances upstream much faster, which is evident from an overprediction in the burning rate and generalized flame surface density. Furthermore, there seem to be smaller cellular structures present, which could be attributed to the small unstable wavelengths that followed from the linear stability analysis. In contrast, including the H-radical diffusion flux in the preferential diffusion of enthalpy results in flame shape, dynamics, and burning rate predictions that closely match the DC benchmark. It is shown that not including the H-radical explicitly leads to a significant underprediction of the tangential enthalpy diffusion flux. The H-radical has a significant tangential component along the flame, and since it carries a significant amount of enthalpy, accurate prediction of enthalpy cross-diffusion requires the explicit inclusion of H-radical diffusion in the enthalpy preferential diffusion term. A final observation is that FGM A also exhibits a slight overprediction in the burning rate, suggesting that enthalpy may play a role as a control variable in accurately capturing thermo-diffusive instabilities. However, further investigation is required to draw definitive conclusions.